OpenStudy (anonymous):

what number decreased by 1/5 of itself equals 84

OpenStudy (anonymous):

Your equation should look something like: x - (x/5)= 84 5x-x = 84(5) 4x = 420 x = 105 Do you understand how I got the equation?

OpenStudy (anonymous):

yes I do thank you

OpenStudy (anonymous):

good! np